The resistance of a wire is 5 ohm at 50°C and 6 ohm at 100°C. The resistance of the wire at 0°C will be ?

Grade:12

2 Answers

Apoorva Arora IIT Roorkee
9 years ago
R=R_{0}(1+\alpha (T-T_{0}))
so
5=R_{0}(1+\alpha (50))
and
6=R_{0}(1+\alpha (100))
dividing,
\alpha =1/200
so
R_{0}= 4 ohms.
Thanks and Regards
Apoorva Arora
IIT Roorkee
askIITians Faculty
Kushagra Madhukar
3 years ago
Dear student,
Please find the solution to your problem.
 
R = Ro(1 + a(T – To))
so
5 = Ro(1 + a(50))
and
6 = Ro(1 + a(100))
dividing,
a = 1/200
so
Ro = 4 ohms
 
Thanks and regards,
Kushagra